I evidently have an opto problem on my MB:
When I turn the game on, I get a Test Report to check switches 74 (pos 5), switch 77 (pos 2), and switch 78 (pos 1).
After I run Drac Test (in T-19), the DMD reads that Drac Pos 1 SW is broken--however, running a Single Switch test in T-3, reveals that 4 of the 5 switches are working perfectly: namely, switches 74 (pos 5), 76 (pos 3), 77 (pos 2), and 78 (pos 1)--they all open and close perfectly when I test them with a tool.
However, I have found a problem with Switch 75 (pos 4): it is currently staying permanently closed.
Earlier (in a T-3 test of Switch 75; pos 4), the switch was continuously cycling on its own from being closed to being open.
I thought it may be a connector problem, so I re-seated all the wires with an IDC insertion tool at the IDC connector to the PCB under the Drac Trac (Defender Switch PCB Assembly A-21402) but that didn't help. On a long-shot, I also re-inserted all the wires on the IDCs leading to the 16-Opto PCB to no avail.
So what's going on here?
A bad opto at U2 on PCB assembly A-21402? A bad transistor on PCB assembly A-21402?
I haven't yet inspected all the transistor pins with a magnifier to check for a cracked pin--I thought I'd stop and ask for help at this point.
Many thanks in advance.